OpenStudy (anonymous):

Can someone help me please. I'm desperate I can get halfway through but I can't figure it out 100%.. Find the x-intercepts of the parabola with vertex (4,-1) and y-intercept (0,15). Answer in this form: (x1,y1),(x2,y2)

OpenStudy (anonymous):

vertex is \((4,-1)\) so you know it looks like \[y=a(x-4)^2-1\] and we can solve for \(a\) because we know that if \(x=0\) then \(y=15\) so set \[15=a(0-4)^2-1\] \[15=16a-1\] \[16a=16\] \[a=1\]

OpenStudy (anonymous):

so your parabola has the equation \[y=(x-4)^2-1\] now you can set it equal to zero and solve

OpenStudy (anonymous):

\[(x-4)^2-1=0\] \[(x-4)^2=1\] \[x-4=1\] or \[x-4=-1\] \[x=5\] or \[x=3\]

OpenStudy (anonymous):

Trying it right now

OpenStudy (anonymous):

OH MY GOODNESS THANK YOU SO MUCH! <3